OpenStudy (anonymous):

3x-2y+2z=30 -x+3y-4z= -33 2x-4y+3z=42

OpenStudy (thcsbabygirl1224):

Ok are doing systems of equations

OpenStudy (thcsbabygirl1224):

First 3x-2y+2z=30 -3x+9y-12z=99 +--------------------- add both equations together

OpenStudy (anonymous):

7y-10z= 129 But if you take 3 times - 33 you get -99. How did you get +99?

OpenStudy (thcsbabygirl1224):

You multiplied -x+3y+4z=-33 by 3